OpenStudy (anonymous):

Jessica is converting 60 feet per hour into yards per hour. Which conversion factor should she use? A. 60 min/1 hour, B. 1 hour/60 min, C. 3 ft/1 yd, D. 1 yd/3 ft

OpenStudy (texaschic101):

3 ft = 1 yd (so 60 ft = 60/3 = 20 yds)

OpenStudy (anonymous):

But which one is that? @texaschic101 Thank you so much

OpenStudy (texaschic101):

it is C

OpenStudy (anonymous):

Thank you!!!! I have some more, would you mind helping?

OpenStudy (texaschic101):

lets see what you got

OpenStudy (anonymous):

Imelda needs to convert 440 centimeters per minute to meters per hour. Which conversion factors should she use? A. 1 cm/0.01 m and 1 h/60 min, B. 1 cm/0.01 m and 60 min/1 h, C. 0.01 m/1 cm and 1 h/60 min, D. 0.01 m/1 cm and 60 min/1 h.

OpenStudy (anonymous):

I think it's C

OpenStudy (texaschic101):

I think it is B.

OpenStudy (anonymous):

What is 84 feet per hour converted into inches per minute? A. 7 inches per minute B. 16.8 inches per minute C. 420 inches per minute D. 5040 inches per minute

OpenStudy (anonymous):

Sorry about the spacing.

OpenStudy (anonymous):

I think it's B, because 12(84)=1008 and 1008/60=16.8

OpenStudy (texaschic101):

84 ft per hour to inches per minute 12 in per foot so 84 ft = 12 x 84 = 1008 60 min = 1 hr 1008/60 = 16.8 I agree with you

OpenStudy (anonymous):

Okay, two more.

OpenStudy (anonymous):

Clara is climbing up a cliff at a rate of about 8 feet per minute. What is her rate in meters per hour? A. about 1574 meters per hour B. about 146 meters per hour C. about 18.3 meters per hour D. about 2.3 meters per hour

OpenStudy (texaschic101):

8 ft per minute to meters per hour it is about 146 meters per hour

OpenStudy (anonymous):

Thanks! Last one: (Assuming I pass. If I don't I have to do it again with similar but different questions) What is 7 miles per hour converted into meters per minute? A. about 13790 meters per minute B. about 188 meters per minute C. about 72.5 meters per minute D. about 26.8 meters per minute

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Nvm I think it's B

OpenStudy (anonymous):

The first two were wrong, so I got 60 and didn't pass and now I have to do it again.

OpenStudy (anonymous):

So I did it again keeping in mind my mistakes and got 100!